OpenStudy (anonymous):

put 14%,1/5, 0.03 in oder from least to greatest

OpenStudy (anonymous):

i will GIVE A MEADEL

OpenStudy (chosenmatt):

turn 1/5 into a percent

OpenStudy (igreen):

14% = 14/100 = 0.14 1/5 = 0.20 So can you order them from least to greatest now? @noterdame12345

OpenStudy (chosenmatt):

then turn 0.03 into a percent then choose least from greatest

OpenStudy (anonymous):

0.003,0.14, 0.20

OpenStudy (chosenmatt):

GOOD JOB!

OpenStudy (anonymous):

THANKS HOW DO YOU GIVE MEADELS

OpenStudy (chosenmatt):

click best response

OpenStudy (chosenmatt):

thanks

OpenStudy (anonymous):

U WELOME